Zn reacts with aqueous CuSO4 according to the equation: Zn +CuSO4 = Cu + ZnSO4
If 100 cm of 0.5mol/dm3 CuSO4 is placed in a calorimeter with a temperature of 18 and enough powdered Zn is added to react with all the Cu ions
The temperature is recorded at 12
Assume that the specific heat capacity of the mixture in the beaker is equal to that of 100 of water.
Calculate the heat change of one mole of Cu ions
1
Expert's answer
2020-06-16T15:03:53-0400

Moles of cuso4 =100*.5/100=.05moles

Enthalpy change={mass of water*specific heat capacity*temperature change}

Mass of water=100gms

Specific heat capacity of water=4.2

Temperature change=18-12=6 degrees

∆H=-(100*4.2*6)

=-2520J

Molar heat change for Cu ions=-2520/.05=-50400J/mol

=-50.4kJ/mol